# Small or Large? Zero-Shot or Finetuned? Guiding Language Model Choice for Specialized Applications in Healthcare

### v1

@@ Abstract
This study aims to guide language model selection by investigating: 1) the necessity of finetuning versus zero-shot usage, 2) the benefits of domain-adjacent versus generic pretrained models, 3) the value of further domain-specific pretraining, and 4) the continued relevance of Small Language Models (SLMs) compared to Large Language Models (LLMs) for specific tasks. Using electronic pathology reports from the British Columbia Cancer Registry (BCCR), three classification scenarios with varying difficulty and data size are evaluated. Models include various SLMs and an LLM. SLMs are evaluated both zero-shot and finetuned; the LLM is evaluated zero-shot only. Finetuning significantly improved SLM performance across all scenarios compared to their zero-shot results. The zero-shot LLM outperformed zero-shot SLMs but was consistently outperformed by finetuned SLMs. Domain-adjacent SLMs generally performed better than the generic SLM after finetuning, especially on harder tasks. Further domain-specific pretraining yielded modest gains on easier tasks but significant improvements on the complex, data-scarce task. The results highlight the critical role of finetuning for SLMs in specialized domains, enabling them to surpass zero-shot LLM performance on targeted classification tasks. Pretraining on domain-adjacent or domain-specific data provides further advantages, particularly for complex problems or limited finetuning data. While LLMs offer strong zero-shot capabilities, their performance on these specific tasks did not match that of appropriately finetuned SLMs. In the era of LLMs, SLMs remain relevant and effective, offering a potentially superior performance-resource trade-off compared to LLMs.
@@ 1 Introduction
The rapid advancement and proliferation of language models (LMs) have revolutionized the field of Natural Language Processing (NLP) [1]. From compact, small language models (SLMs) like BERT [2] to massive large language models (LLMs) like GPT-4, practitioners now have more choices than ever. This raises a critical question: in the era of language models, which LM should be used for a given task? The decision involves several considerations.
@@
    With the rise of LLMs, are SLMs like BERT still relevant?
@@ 2 Methods and Metrics
For empirical evaluation, we use several off-the-shelf SLMs: a strong general-purpose model like RoBERTa (125M parameters) [5], and models specifically pre-trained on broader clinical text (pathologyBERT (108M parameters) [6], and Gatortron (345M parameters) [7]). For our LLM, we use Mistral nemo instruct (12B parameters) [^1] https://huggingface.co/nvidia/Mistral-NeMo-12B-Instruct [8] selected as a representative high-performance open-source LLM. We further pretrain RoBERTa and Gatortron on 1M pathology reports from BCCR using Masked Language Modeling (MLM) and denote these domain specific pretrained models as BCCRoBERTa and BCCRTron. For performance reporting, we use the macro-averaged F1 scores.
@@
For empirical evaluation using our scenarios, we observe that for SLMs, finetuning consistently significantly outperforms the zero-shot approaches where zero-shot performance ranges from 0.34-0.40 for scenario (a), 0.01 for scenario (b), and 0.02-0.13 for scenario (c). After finetuning, the performance increased to 0.95-0.97 for scenario (a), 0.78-0.85 for scenario (b), and 0.60-0.78 for scenario (c). For LLM, in its zero-shot capacity, it reached the performance of 0.76 for scenario (a), 0.54 for scenario (b), and 0.65 for scenario (c). LLM zero-shot performance was evaluated using task-specific instructions provided via prompting. To mirror resource constraints encountered in the real-world applications, we do not finetune the LLM. We provide detailed results in a Table 1.
@@ 5 Revisiting the Questions
With the help of the information provided above, we conclude the paper by revisiting the questions and providing our recommendations based on our experience and empirical evidence.
@@
    Do we need to finetune a model or can we use it in its zero-shot capacity? Our evaluation clearly shows that finetuned models significantly outperform their zero-shot counterparts, especially for SLMs, as they have limited zero-shot capabilities. LLMs, as anticipated perform better in zero-shot capacity than their SLM counterparts, but are significantly outperformed by finetuned SLMs.
    Our recommendation: SLMs should be finetuned on the task-specific dataset. A finetuned SLM outperforms a zero-shot LLM for many well-defined tasks, especially where the models are intended for use in a specialized domain (cancer pathology reports in our case).
@@
    Our recommendation: If the data and compute resources are available, it is worthwhile further pretraining SLMs for extra performance boost.
@@
    With the rise of LLMs, are SLMs like BERT still relevant? We have shown that for specialized tasks, SLMs still endure and outperform zero-shot LLMs. While LLMs excel at generation, complex reasoning, and few-shot learning across a vast range of topics, SLMs often provide sufficient or even superior performance for many well-defined NLP tasks, such as text classification, named entity recognition,etc. particularly when task-specific training data is available for finetuning. Further, for resource-constrained environments, which are a common place, SLMs offer a compelling balance of performance and efficiency.
    Our recommendation: If the task is well defined (example: classification) and the data is available for finetuning, SLMs should be the first choice.

### v2

@@ Abstract
This study aims to guide language model selection by investigating: 1) the necessity of finetuning versus zero-shot usage, 2) the benefits of domain-adjacent versus generic pretrained models, 3) the value of further domain-specific pretraining, and 4) the continued relevance of BERT-style, bidirectional Language Models (BiLMs) compared to modern Large Language Models (LLMs) for specific tasks. Using electronic pathology reports from the British Columbia Cancer Registry (BCCR), three classification scenarios with varying difficulty and data size are evaluated. Models include various BiLMs and an LLM. BiLMs are evaluated both zero-shot and finetuned; the LLM is evaluated zero-shot only. Finetuning significantly improved BiLMs performance across all scenarios compared to their zero-shot results. The zero-shot LLM outperformed zero-shot BiLMs but was consistently outperformed by finetuned BiLMs. Domain-adjacent BiLMs generally performed better than the generic BiLM after finetuning, especially on harder tasks. Further domain-specific pretraining yielded modest gains on easier tasks but significant improvements on the complex, data-scarce task. The results highlight the critical role of finetuning for BiLMs in specialized domains, enabling them to surpass zero-shot LLM performance on targeted classification tasks. Pretraining on domain-adjacent or domain-specific data provides further advantages, particularly for complex problems or limited finetuning data. While LLMs offer strong zero-shot capabilities, their performance on these specific tasks did not match that of appropriately finetuned BiLMs. In the era of LLMs, BiLMs remain relevant and effective, offering a potentially superior performance-resource trade-off compared to LLMs.
@@ 1 Introduction
The rapid advancement and proliferation of language models (LMs) have revolutionized the field of Natural Language Processing (NLP) [1]. From compact, BERT-style, bidirectional language models (BiLMs) [2] to modern, massive large language models (LLMs) like GPT-4, practitioners now have more choices than ever. This raises a critical question: in the era of language models, which LM should be used for a given task? The decision involves several considerations.
@@
    With the rise of LLMs, are BiLMs like BERT still relevant?
@@ 2 Methods and Metrics
For empirical evaluation, we use several off-the-shelf BiLMs: a strong general-purpose model like RoBERTa (125M parameters) [5], and models specifically pre-trained on broader clinical text (pathologyBERT (108M parameters) [6], and Gatortron (345M parameters) [7]). For our LLM, we use Mistral nemo instruct (12B parameters) [^1] https://huggingface.co/nvidia/Mistral-NeMo-12B-Instruct [8] selected as a representative high-performance open-source LLM. We further pretrain RoBERTa and Gatortron on 1M pathology reports from BCCR using Masked Language Modeling (MLM) and denote these domain specific pretrained models as BCCRoBERTa and BCCRTron. For performance reporting, we use the macro-averaged F1 scores.
@@
For empirical evaluation using our scenarios, we observe that for BiLMs, finetuning consistently significantly outperforms the zero-shot approaches where zero-shot performance ranges from 0.34-0.40 for scenario (a), 0.01 for scenario (b), and 0.02-0.13 for scenario (c). After finetuning, the performance increased to 0.95-0.97 for scenario (a), 0.78-0.85 for scenario (b), and 0.60-0.78 for scenario (c). For LLM, in its zero-shot capacity, it reached the performance of 0.76 for scenario (a), 0.54 for scenario (b), and 0.65 for scenario (c). LLM zero-shot performance was evaluated using task-specific instructions provided via prompting. To mirror resource constraints encountered in the real-world applications, we do not finetune the LLM. We provide detailed results in a Table 1. The results align with expectations, as BiLMs lack the zero-shot capabilities of modern LLMs, particularly for out-of-domain tasks.
@@ 5 Revisiting the Questions
With the help of the information provided above, we conclude the paper by revisiting the questions and providing our recommendations based on our experience and empirical evidence, while acknowledging the limitation that we have only tested a small set of models.
@@
    Do we need to finetune a model or can we use it in its zero-shot capacity? Our evaluation clearly shows that finetuned models significantly outperform their zero-shot counterparts, especially for BiLMs, as they have limited zero-shot capabilities. LLMs, as anticipated perform better in zero-shot capacity than their BiLM counterparts, but are significantly outperformed by finetuned BiLMs.
    Our recommendation: BiLMs should be finetuned on the task-specific dataset. A finetuned BiLM outperforms a zero-shot LLM for many well-defined tasks, especially where the models are intended for use in a specialized domain (cancer pathology reports in our case).
@@
    Our recommendation: If the data and compute resources are available, it is worthwhile further pretraining BiLMs for extra performance boost.
@@
    With the rise of LLMs, are BiLMs like BERT still relevant? We have shown that for specialized tasks, BiLMs still endure and outperform zero-shot LLMs. While LLMs excel at generation, complex reasoning, and few-shot learning across a vast range of topics, BiLMs often provide sufficient or even superior performance for many well-defined NLP tasks, such as text classification, named entity recognition,etc. particularly when task-specific training data is available for finetuning. Further, for resource-constrained environments, which are a common place, BiLMs offer a compelling balance of performance and efficiency.
    Our recommendation: If the task is well defined (example: classification) and the data is available for finetuning, BiLMs should be the first choice.